# Oriented matroids and type $\mathbb{A}$ cluster categories

### v2

@@ 4.3. Triangulation to maximal green sequence
We now describe the inverse construction, which starts with a stackable triangulation $\Delta$ of $\mathcal{M}_{\mathsf{T}}$, and outputs an equivalence class of maximal green sequences $\mathcal{G}]$ of $\mathsf{T}$. As in the previous section, we have fixed a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, and have let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$. Given a triangulation $T_{i}$ of the $m$-gon which contains a quadrilateral abcdabcd, then flipping the diagonal in abcdabcd is called the increasing flip given by abcdabcd if it replaces the lower facets of abcdabcd with respect to $T_{i}$ by the upper facets of abcdabcd with respect to $T_{i}$. If the reverse is true, then it is called a decreasing flip.
@@
For the base case, $i=1$, let $F$ be a lower facet of $\sigma_{1}$. By Definition 2.5(2), we have that $F$ is either a facet of $\mathcal{M}_{T}$, or is contained in precisely two simplices of $\Delta$. If $F$ is contained in another simplex $\sigma_{j}$ of $\Delta$, then it must be an upper facet of $\sigma_{j}$ by Lemma 3.14. However, then we would have $\sigma_{j}\rightsquigarrow\sigma_{1}$, which would imply that $\sigma_{j}<\sigma_{1}$, since $\leqslant$ is a linear extension of $\rightsquigarrow$. Hence, all lower facets of $\sigma_{1}$ are facets of $\mathcal{M}_{T}$, and hence lower facets of $\mathcal{M}_{T}$. Since the lower facets of $\mathcal{M}_{T}$ are given by $T+1$, the base case holds.
@@
Now suppose that $F$ is contained in precisely one other simplex $\sigma_{j}$ of $\Delta$. By Lemma 3.14, we have that $F$ must be an upper facet of $\sigma_{j}$, and so we have $\sigma_{j}\leqslant\sigma_{i}$. By the induction hypothesis, we have that the lower facets of $\sigma_{j}$ are contained in $T_{j-1}$, and so $T_{j}$ contains the upper facets of $\sigma_{j}$, and so contains $F$. Since $F$ is only contained in $\sigma_{j}$ and $\sigma_{i}$ by Definition 2.5(2), it remains a triangle of $T_{i-1}$.
@@
Construction 4.9.
Let $\Delta$ be a stackable triangulation of $\mathcal{M}_{T}$. As before, since $\rightsquigarrow$ is a partial order on $\Delta$, we may take a linear extension of it to a total order $\leqslant$. Let $\sigma_{1},\sigma_{2},\dots,\sigma_{l}$ be the simplices of $\Delta$, ordered according to $\leqslant$.
We then take the sequence $T_{0}$, $T_{1}$, …, $T_{l}$ of triangulations from Proposition 4.8. By Section 2.2 and Lemma 4.1, this corresponds to a sequence of cluster-tilting objects $\mathsf{T}_{0}$, $\mathsf{T}_{1}$, …, $\mathsf{T}_{l}$ of $\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{i}$ is a green mutation of $\mathsf{T}_{i-1}$, and $\mathsf{T}_{l}=\Sigma\mathsf{T}$. Thus this gives a maximal green sequence $\mathcal{G}^{\leqslant}_{\Delta}$ of $\mathsf{T}$, and hence an equivalence class $\mathcal{G}^{\leqslant}_{\Delta}]$.

### v3

@@ 4.3. Triangulation to maximal green sequence
We now describe the inverse construction, which starts with a stackable triangulation $\Delta$ of $\mathcal{M}_{\mathsf{T}}$, and outputs an equivalence class of maximal green sequences $\mathcal{G}]$ of $\mathsf{T}$. As in the previous section, we have fixed a cluster-tilting object $\mathsf{T}$ in $\mathscr{C}_{n}$, and have let $T$ be the triangulation of the $(n+3)$-gon corresponding to $\Sigma\mathsf{T}$.
@@
For the base case, $i=1$, let $F$ be a lower facet of $\sigma_{1}$. By Definition 2.6(1), we have that $F$ is either a facet of $\mathcal{M}_{T}$, or is contained in precisely two simplices of $\Delta$. If $F$ is contained in another simplex $\sigma_{j}$ of $\Delta$, then it must be an upper facet of $\sigma_{j}$ by Lemma 3.15. However, then we would have $\sigma_{j}\rightsquigarrow\sigma_{1}$, which would imply that $\sigma_{j}<\sigma_{1}$, since $\leqslant$ is a linear extension of $\rightsquigarrow$. Hence, all lower facets of $\sigma_{1}$ are facets of $\mathcal{M}_{T}$. It is clear from the definition of lower facets that they are in fact lower facets of $\mathcal{M}_{T}$. Since the lower facets of $\mathcal{M}_{T}$ are given by $T+1$, the base case holds.
@@
Now suppose that $F$ is contained in precisely one other simplex $\sigma_{j}$ of $\Delta$. By Lemma 3.15, we have that $F$ must be an upper facet of $\sigma_{j}$, and so we have $\sigma_{j}\leqslant\sigma_{i}$. By the induction hypothesis, we have that the lower facets of $\sigma_{j}$ are contained in $T_{j-1}$, and so $T_{j}$ contains the upper facets of $\sigma_{j}$, and so contains $F$. Since $F$ is only contained in $\sigma_{j}$ and $\sigma_{i}$ by Definition 2.6(1), it must still be a triangle of $T_{i-1}$.
@@
Construction 4.11.
Let $\Delta$ be a stackable triangulation of $\mathcal{M}_{T}$. As before, since $\rightsquigarrow$ is a partial order on $\Delta$, we may extend it to a total order $\leqslant$. Let $\sigma_{1},\sigma_{2},\dots,\sigma_{l}$ be the simplices of $\Delta$, ordered according to $\leqslant$.
We then take the sequence $T_{0}$, $T_{1}$, …, $T_{l}$ of triangulations from Proposition 4.10. By Section 2.2 and Lemma 4.1, this corresponds to a sequence of cluster-tilting objects $\mathsf{T}_{0}$, $\mathsf{T}_{1}$, …, $\mathsf{T}_{l}$ of $\mathscr{C}_{n}$ such that $\mathsf{T}_{0}=\mathsf{T}$, $\mathsf{T}_{i}$ is a green mutation of $\mathsf{T}_{i-1}$, and $\mathsf{T}_{l}=\Sigma\mathsf{T}$. This therefore gives a maximal green sequence $\mathcal{G}^{\leqslant}_{\Delta}$ of $\mathsf{T}$, and hence an equivalence class $\mathcal{G}^{\leqslant}_{\Delta}]$.
@@
Appendix A Module, derived, and cluster categories of type $\mathbb{A}$
In this appendix, we prove Lemma 4.2 using combinatorial models for the representation theory of the path algebra of linearly oriented type $\mathbb{A}$. Our aim is partly expository, for readers less familiar with quiver representations. We denote the quiver
$1\leftarrow 2\leftarrow\dots\leftarrow n$
by $\mathbb{A}_{n}$ and its path algebra over a field $K$ by $K\mathbb{A}_{n}$.
A.1. The module category
For background on quiver representations we refer to [Sch14] or [ASS06]. It is well-known that finite-dimensional right $KQ$-modules are equivalent to finite-dimensional representations of $Q$ over $K$ [ASS06, Chapter III, Theorem 1.6]. It is also well-known that the indecomposable representations of $\mathbb{A}_{n}$ are given by
$0\to\dots\to\overset{i}{K}\xleftarrow{\operatorname{id}_{K}}\overset{i+1}{K}\xleftarrow{\operatorname{id}_{K}}\dots\xleftarrow{\operatorname{id}_{K}}\overset{j}{K}\to 0\to\dots\to 0,$
where $1\leqslant i\leqslant j\leqslant n$. This follows for instance from Gabriel’s Theorem [Gab72]. We denote this indecomposable representation by $\mathsf{M}_{i,j+2}$. Hence, the indecomposable representations of $\mathbb{A}_{n}$ are $\{\mathsf{M}_{ac}:a,c\in[n+2]\text{ with }a+2\leqslant c\}$.
Suppose we have a homomorphism of indecomposable representations representations $f\colon\mathsf{M}_{ac}\to\mathsf{M}_{bd}$ given by
(A.1) $V_{1}$$V_{2}$$\dots$$V_{n}$$W_{1}$$W_{2}$$\dots$$W_{n},$$\scriptstyle{f_{1}}$$\scriptstyle{g_{1}}$$\scriptstyle{f_{2}}$$\scriptstyle{g_{2}}$$\scriptstyle{f_{n}}$$\scriptstyle{g_{n-1}}$$\scriptstyle{h_{1}}$$\scriptstyle{h_{2}}$$\scriptstyle{h_{n-1}}$
where $\mathsf{M}_{ac}$ is $V_{1}\leftarrow\dots\leftarrow V_{n}$ and $\mathsf{M}_{bd}$ is $W_{1}\leftarrow\dots\leftarrow W_{n}$. If a1<b<c1<da-1<b<c-1<d, there is a non-zero homomorphism $f_{ac,bd}\colon\mathsf{M}_{ac}\to\mathsf{M}_{bd}$ which factorises as
$\mathsf{M}_{ac}\twoheadrightarrow\mathsf{M}_{bc}\hookrightarrow\mathsf{M}_{bd}.$
This homomorphism is in fact unique up to scalar, since the choice of scalar for one of the non-zero components $f_{i}$ for $i\in[b,c]$ determines all of the others, since the squares in the diagram commute. By inspecting the diagram (A.1), note further that $\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{ac},\mathsf{M}_{bd})=0$ if any of the following hold: $b\leqslant a-1$, $c\leqslant b+1$, $d\leqslant c-1$. We have therefore shown the following lemma.
Lemma A.1 ([OT12, Theorem 3.6(3)]).
We have that $\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{ac},\mathsf{M}_{bd})\not\cong 0$ if and only if a1<b<c1<da-1<b<c-1<d, and in this case $\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{ac},\mathsf{M}_{bd})\cong K$.
We can go further and give a criterion for when two non-zero morphisms compose to give a non-zero morphisms. (Experts will note that this is just another way of rewriting the well-known relations for the Auslander algebra.)
Lemma A.2.
A pair of non-zero morphisms $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{q_{0}q_{1}}$ and $\mathsf{M}_{q_{0}q_{1}}\to\mathsf{M}_{r_{0}r_{1}}$ compose to give a non-zero morphism $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{r_{0}r_{1}}$ if and only if $p_{0}-1\leqslant q_{0}-1<r_{0}<p_{1}-1\leqslant q_{1}-1<r_{1}$.
Proof.
We have that $p_{0}-1\leqslant q_{0}-1<r_{0}<p_{1}-1\leqslant q_{1}-1<r_{1}$ if and only if the following three sets of inequalities all hold.
$p_{0}-1<q_{0}<p_{1}-1<q_{1},$
$q_{0}-1<r_{0}<q_{1}-1<r_{1},$
$p_{0}-1<r_{0}<p_{1}-1<r_{1}.$
By Lemma A.1, we have that these three sets of inequalities are equivalent to the non-vanishing of the Hom-spaces
$\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{p_{0}p_{1}},\mathsf{M}_{q_{0}q_{1}})$ $\neq 0,$
$\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{q_{0}q_{1}},\mathsf{M}_{r_{0}r_{1}})$ $\neq 0,$
$\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{p_{0}p_{1}},\mathsf{M}_{r_{0}r_{1}})$ $\neq 0.$
Hence, the lemma is equivalent to the statement that two morphisms $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{q_{0}q_{1}}$ and $\mathsf{M}_{q_{0}q_{1}}\to\mathsf{M}_{r_{0}r_{1}}$ to give a non-zero morphism $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{r_{0}r_{1}}$ if and only if there exists a non-zero morphism $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{r_{0}r_{1}}$. The forwards direction is immediate. For the backwards direction, suppose that we have two non-zero morphisms $\mathsf{M}_{p_{0}p_{1}}\to\mathsf{M}_{q_{0}q_{1}}$ and $\mathsf{M}_{q_{0}q_{1}}\to\mathsf{M}_{r_{0}r_{1}}$. By Lemma A.1, these morphisms are unique up to scalar, so we may assume without loss of generality that these morphisms are $f_{p_{0}p_{1},q_{0}q_{1}}$ and $f_{q_{0}q_{1},r_{0}r_{1}}$. Then suppose that we have that $\operatorname{Hom}_{K\mathbb{A}_{n}}(\mathsf{M}_{p_{0}p_{1}},\mathsf{M}_{r_{0}r_{1}})\neq 0$, so that in particular we have the non-zero morphism $f_{p_{0}p_{1},r_{0}r_{1}}$.
It suffices to show that $f_{q_{0}q_{1},r_{0}r_{1}}\circ f_{p_{0}p_{1},q_{0}q_{1}}=f_{p_{0}p_{1},r_{0}r_{1}}$. This follows from the fact that the diagram
$\mathsf{M}_{p_{0}p_{1}}$$\mathsf{M}_{q_{0}q_{1}}$$\mathsf{M}_{r_{0}r_{1}}$$\mathsf{M}_{q_{0}p_{1}}$$\mathsf{M}_{r_{0}q_{1}}$$\mathsf{M}_{r_{0}p_{1}}$}
commutes. Indeed, $f_{p_{0}p_{1},q_{0}q_{1}}$ is the composition $\mathsf{M}_{p_{0}p_{1}}\twoheadrightarrow\mathsf{M}_{q_{0}p_{1}}\hookrightarrow\mathsf{M}_{q_{0}q_{1}}$; $f_{q_{0}q_{1},r_{0}r_{1}}$ is the composition $\mathsf{M}_{q_{0}q_{1}}\twoheadrightarrow\mathsf{M}_{r_{0}q_{1}}\hookrightarrow\mathsf{M}_{r_{0}r_{1}}$; and $f_{p_{0}p_{1},r_{0}r_{1}}$ is the composition $\mathsf{M}_{p_{0}p_{1}}\twoheadrightarrow\mathsf{M}_{q_{0}p_{1}}\twoheadrightarrow\mathsf{M}_{r_{0}p_{1}}\hookrightarrow\mathsf{M}_{r_{0}q_{1}}\hookrightarrow\mathsf{M}_{r_{0}r_{1}}$. Commutativity of the middle square follows easily from considering the explicit quiver representations: the morphism given by either path around the square is precisely the map which is zero on the vector spaces at vertices in $q_{0},r_{0}-1]$ and the identity on the vector spaces at vertices in $r_{0},p_{1}-2]$. ∎
A.2. The derived category
We now describe the bounded derived category $\mathscr{D}_{n}:=\mathscr{D}^{b}(K\mathbb{A}_{n})$ of right $K\mathbb{A}_{n}$-modules. For $a,c\in[n+2]$ with $a+2\leqslant c$, we define $\mathsf{U}_{ac}$ to be the complex given by $\mathsf{M}_{ac}$ concentrated in degree zero. We then define
(A.2) $\mathsf{U}_{ac}[1]$ $=\mathsf{U}_{c-1,a+n+2},$
(A.3) $\mathsf{U}_{ac}[-1]$ $=\mathsf{U}_{c-(n+2),a+1}.$
Hence, every complex which is given by an indecomposable $K\mathbb{A}_{n}$-module concentrated in one degree is labelled as $\mathsf{U}_{ac}$ for some $a,c$. By a result of Happel [Hap88, Lemma 5.2], indecomposable objects in $\mathscr{D}_{n}$ are precisely given by complexes of indecomposable $K\mathbb{A}_{n}$-modules concentrated in one degree. Hence, the labelling of the indecomposable objects $\mathsf{U}_{ac}$ gives a bijection between the indecomposable objects of $\mathscr{D}_{n}$ and the elements of $\{(a,c)\in\mathbb{Z}^{2}:a+2\leqslant c\leqslant a+n+1\}$.
This labelling is a special case of the labelling in [OT12, Section 6], as follows from combining [OT12, Lemma 6.6] and [OT12, Lemma 6.7]. We then have the following analogue of Lemma A.1 for $\mathscr{D}_{n}$.
Lemma A.3 ([OT12, Proof of Proposition 6.1]).
We have $\operatorname{Hom}_{\mathscr{D}_{n}}(\mathsf{U}_{ac},\mathsf{U}_{bd})\not\cong 0$ if and only if a1<b<c1<d<a+n+2a-1<b<c-1<d<a+n+2, and in this case $\operatorname{Hom}_{\mathscr{D}_{n}}(\mathsf{U}_{ac},\mathsf{U}_{bd})\cong K$.
Recalling the Nakayama functor $\nu$ from Section 2.2, we denote by $\tau$ the functor $\nu[-1]$, which is an auto-equivalence of $\mathscr{D}_{n}$ known as the Auslander–Reiten translate. It admits the following combinatorial description.
Lemma A.4 ([OT12, Lemma 6.6]).
We have that $\tau\mathsf{U}_{ac}=\mathsf{U}_{a-1,c-1}$.
We can then use this to prove the following analogue of Lemma A.2.
Lemma A.5.
A pair of non-zero morphisms $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{q_{0}q_{1}}$ and $\mathsf{U}_{q_{0}q_{1}}\to\mathsf{U}_{r_{0}r_{1}}$ compose to give a non-zero morphism $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{r_{0}r_{1}}$ if and only if $p_{0}-1\leqslant q_{0}-1<r_{0}<p_{1}-1\leqslant q_{1}-1<r_{1}<p_{0}+n+2$.
Proof.
Suppose that we have $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{q_{0}q_{1}}$ and $\mathsf{U}_{q_{0}q_{1}}\to\mathsf{U}_{r_{0}r_{1}}$ composing to give a non-zero morphism $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{r_{0}r_{1}}$. Then, by applying the auto-equivalence $\tau^{p_{0}-1}$, we can assume that $p_{0}=1$, so that $\mathsf{U}_{p_{0}p_{1}}$ is concentrated in degree zero. Then the existence of non-zero morphisms $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{q_{0}q_{1}}$ gives by Lemma A.3 that $p_{0}\leqslant q_{0}$ and $q_{1}\leqslant p_{0}+n+1=n+2$, so that $\mathsf{U}_{q_{0}q_{1}}$ is concentrated in degree zero too. We likewise deduce that $\mathsf{U}_{r_{0}r_{1}}$ is concentrated in degree zero. We then use Lemma A.2 to deduce that $p_{0}-1\leqslant q_{0}-1<r_{0}<p_{1}-1\leqslant q_{1}-1<r_{1}$. The fact that $r_{1}<n+3=p_{0}+n+2$ follows from the fact that $\mathsf{U}_{r_{0}r_{1}}$ is concentrated in degree zero.
Conversely, suppose that $p_{0}-1\leqslant q_{0}-1<r_{0}<p_{1}-1\leqslant q_{1}-1<r_{1}<p_{0}+n+2$. Then we may again apply $\tau^{p_{0}-1}$, whence these inequalities give us that all of $\mathsf{U}_{p_{0}p_{1}}$, $\mathsf{U}_{q_{0}q_{1}}$, and $\mathsf{U}_{r_{0}r_{1}}$ are concentrated in degree zero. We can then apply Lemma A.2 to deduce that we have a pair of non-zero morphisms $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{q_{0}q_{1}}$ and $\mathsf{U}_{q_{0}q_{1}}\to\mathsf{U}_{r_{0}r_{1}}$ composing to give a non-zero morphism $\mathsf{U}_{p_{0}p_{1}}\to\mathsf{U}_{r_{0}r_{1}}$. ∎
A.3. The cluster category
We now deduce similar results for the cluster category $\mathscr{C}_{n}$. The following description of $\nu[-2]$ can be obtained from combining (A.3) with Lemma A.4.
Lemma A.6 ([OT12, Lemma 6.7]).
We have that $\nu\mathsf{U}_{ac}[-2]=\mathsf{U}_{c-(n+3),a}$.
Let $\psi\colon\mathscr{D}_{n}\to\mathscr{C}_{n}$ be the natural projection functor sending an object to its orbit. Lemma A.6 then allows us to label the indecomposable objects of the $\mathscr{C}_{n}$ as $\mathsf{O}_{ac}$ such that
(A.4) $\psi(\mathsf{U}_{a^{\prime}c^{\prime}})=\mathsf{O}_{ac}\text{ if and only if }\{a^{\prime},c^{\prime}\}\equiv\{a,c\}\mod(n+3).$
The condition for $\operatorname{Hom}_{\mathscr{C}_{n}}(\mathsf{O}_{ac},\mathsf{O}_{bd})\not\cong 0$ from Section 2.2 ($a-1\prec b\prec c-1\prec d$) then follows from (A.4) with Lemma A.3. As always, we have that the non-zero morphism is unique up to scalar if it exists.
We can now prove Lemma 4.2, which is the analogue of Lemma A.2 and Lemma A.5 for the cluster category.
Proof of Lemma 4.2.
Suppose that $\mathsf{O}_{P}\to\mathsf{O}_{Q}$ and $\mathsf{O}_{Q}\to\mathsf{O}_{R}$ in $\mathscr{C}_{n}$ compose to give a non-zero morphism $\mathsf{O}_{P}\to\mathsf{O}_{R}$. Then, since $\mathscr{C}_{n}$ is an orbit category of $\mathscr{D}_{n}$, we can lift these morphisms to morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ which compose to give a non-zero map, where $\psi(\mathsf{U}_{P^{\prime}})=\mathsf{O}_{P}$, $\psi(\mathsf{U}_{Q^{\prime}})=\mathsf{O}_{Q}$, and $\psi(\mathsf{U}_{R^{\prime}})=\mathsf{O}_{R}$. We have non-zero morphisms $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{Q^{\prime}}$ and $\mathsf{U}_{Q^{\prime}}\to\mathsf{U}_{R^{\prime}}$ in $\mathscr{D}_{n}$ composing to give a non-zero morphism $\mathsf{U}_{P^{\prime}}\to\mathsf{U}_{R^{\prime}}$ if and only if
(A.5) $p^{\prime}_{0}-1\leqslant q^{\prime}_{0}-1<r^{\prime}_{0}<p^{\prime}_{1}-1\leqslant q^{\prime}_{1}-1<r^{\prime}_{1}<p^{\prime}_{0}+n+2$
by Lemma A.5, where $P^{\prime}=p^{\prime}_{0}p^{\prime}_{1}$, $Q^{\prime}=q^{\prime}_{0}q^{\prime}_{1}$, and $R^{\prime}=r^{\prime}_{0}r^{\prime}_{1}$. Applying (A.4) gives us the desired cyclic ordering $p_{0}-1\preccurlyeq q_{0}-1\prec r_{0}\prec p_{1}-1\preccurlyeq q_{1}-1\prec r_{1}$.
Conversely, if such a cyclic ordering holds, one can lift $\mathsf{O}_{P}$, $\mathsf{O}_{Q}$, and $\mathsf{O}_{R}$ to objects $\mathsf{U}_{P^{\prime}}$, $\mathsf{U}_{Q^{\prime}}$, and $\mathsf{U}_{R^{\prime}}$ of $\mathscr{D}_{n}$ such that (A.5) holds. This gives that the morphisms in $\mathscr{D}_{n}$ compose to give a non-zero morphism by Lemma A.5, and so the morphisms in $\mathscr{C}_{n}$ likewise compose to give a non-zero morphism. ∎